OpenStudy (anonymous):

Which of the following represents 3x - 5y + 10 = 0 written in slope-intercept form? y= -x+ 2 y=x- 2 y=x+ 2

mathslover (mathslover):

Is it in options?

OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)? 2x- 3y- 5 = 0 2x+ 3y- 5 = 0 -2x+ 3y- 5 = 0

mathslover (mathslover):

Which question you want to be helped @brittnicoleee_ ?

OpenStudy (anonymous):

What is the equation of the line that passes through (4, -1) and (-2, 3)? 2x- 3y- 5 = 0 2x+ 3y- 5 = 0 -2x+ 3y- 5 = 0